Next year George will be twice as old as Marshall will be then. Five years ago George was 8 times as old as Marshall was then. What are their ages now?
Answer by stanbon(75887)

Next year George will be twice as old as Marshall will be then. Five years ago George was 8 times as old as Marshall was then. What are their ages now?
-----
G+1 = 2(M+1)
G-5 = 8(M-5)
----------------------
Substitute for "G" and solve for "M"::
2M+2-1-5 = 8M-40
6M = 36
Marshall = 6 yrs
G = 2*M+2-1 = 12+2-1 = 13 yrs
----------------------
Cheers,
Stan H.
------------